OpenStudy (anonymous):

The length of one leg of a right triangle is 20 ft and the length of the hypotenuse is 52 ft. What is the length of the other leg? A. 5.7 feet B. 32.0 feet C. 48.0 feet D. 70.8 feet

OpenStudy (anonymous):

r u in 7th grade

OpenStudy (anonymous):

Yes why @iluv2hlp

OpenStudy (anonymous):

do u do k12

OpenStudy (anonymous):

Yup @iluv2hlp

OpenStudy (anonymous):

Can you please help me with my question? @iluv2hlp

OpenStudy (anonymous):

what state

OpenStudy (anonymous):

California

OpenStudy (anonymous):

Can you please help?

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

i do cava

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I do cava too :)

OpenStudy (anonymous):

cool whats ur name

OpenStudy (anonymous):

Use Pythagorean Theorem a^2+b^2=c^2 20^2+b^2=52^2 400+b^2=2704 b^2=2304 b=48 Answer is C.